40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jul 14 16:52:47 2020, gautam, Update, Electronics, Coil drivers for the test masses coilDriverSchem.pdf
    Reply  Wed Jul 15 21:08:43 2020, gautam, Update, Electronics, ETM coil outputs DQed CDS.pngcoilOutDQed.png
    Reply  Mon Jul 20 17:23:46 2020, gautam, Update, Electronics, Coil drivers for the test masses ALSlock_timeDomain.pdfALSlock.pdfRFlock.pdf
Message ID: 15488     Entry time: Wed Jul 15 21:08:43 2020     In reply to: 15480
Author: gautam 
Type: Update 
Category: Electronics 
Subject: ETM coil outputs DQed 

To facilitate this investigation, I've DQed the 4 face coil outputs for the two ETMs. EX is currently running with 5 times the series resistance of EY, so it'll be a nice consistency check. Compilation, installation etc went smooth. But when restarting the c1scx model, there was a weird issue - the foton file, C1SCX.txt, got completely wiped (all filter coefficients were empty, even though the filter module names themselves existed). I just copied the chiara backup version, restarted the model, and all was well again.

This corresponds to 8 additional channels, recorded at 16k as float 32 numbers, so in the worst case (neglecting any clever compression algorithms), we are using disk space at a rate of ~4 MB/s more. Seems okay, but anyway, I will remove these DQ channels in a few days, once we're happy we have enough info to inform the coil driver design.

spoke too soon - there was an RFM error for the TRX channel, and restarting that model on c1sus took down all the vertex FEs. Anyways, now, things are back to normal I think. The remaining red light in c1lsc is from the DNN model not running - I forgot to remove those channels, this would've been a good chance! Anyways, given that there is an MLTI in construction, I'm removing these channels from the c1lsc model, so the next time we restart, the changes will be propagated.

For whatever reason, my usual locking scripts aren't able to get me to the PRFPMI locked state - some EPICS channel value must not have been set correctly after the model reboot 😞. I'll debug in the coming days.

Fun times lie ahead for getting the new BHD FEs installed I guess 🤡 ....

Quote:
 

Looking at signals to the ETMs from the current lock acquisition sequence, the RMS current to a single coil is approximately _____ (to be filled in later).

So we may need a version of the fast coil driver that supports a low noise mode (with large series resistance) and a high-range mode (with lower series resistance for lock acquisition).

Attachment 1: CDS.png  31 kB  Uploaded Wed Jul 15 22:46:07 2020  | Hide | Hide all
CDS.png
Attachment 2: coilOutDQed.png  149 kB  Uploaded Wed Jul 15 22:50:04 2020  | Hide | Hide all
coilOutDQed.png
ELOG V3.1.3-